OpenStudy (anonymous):

Write the solution for the following problem as an ordered pair: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (anonymous):

x + x +y = 46 y = x + 4 x + x + y = 46 x + x +(x + 4) = 46 x +x + x + 4 = 46 3x + 4 = 46 3x = 46 - 4 3x = 42 x = 42/3 x = 14 y = x + 4 y = 14 + 4 y = 20 so since x = 14 and y = 20 it's (14,20) that's the answer